OpenStudy (anonymous):

Please help! A chemist dissolves 1.000 x 10^3 g of potassium hydroxide in enough water to make 1.5 L of stock solution. The chemist later wants 2.00 x 10^2 mL of 0.10 M potassium hydroxide. How much stock solution does the chemist need in order to make this second solution?

OpenStudy (anonymous):

c1V1=c2V2

OpenStudy (anonymous):

simply use M1V1=M2V2 this concept is for this type of questions here v= volume , m= molarity

OpenStudy (anonymous):

first you need to get amount of potassium hydroxide from mass: n=m/M then you calculate concentration c=n/V then you have been given the amount of solution you need and concentration so you just put it in c1V1=c2V2 V1= c2V2/c1

OpenStudy (anonymous):

c1 = 16,66 M so V1= 0,1M * 0,2 dm3 / 16,66M=1,2 * 10^-3 dm3 or 1,2 mL

OpenStudy (anonymous):

this concept is very importannt in day to day lyf.,.

OpenStudy (anonymous):

How did you get C1 = 16,66M?

OpenStudy (anonymous):

c=n/V n=m/M

OpenStudy (anonymous):

So 1000g of KOH = 16,66 M?

OpenStudy (anonymous):

n= 1*10^3g / 40 gmol-1 = 25 mol c= 25 mol/1,5 dm3 = 16,66 M